OpenStudy (anonymous):

sam and nora leave their business office at the same time. Sam drives north at 53 km/h Nora drives south at 47 km/h. How long until they are 50 km apart

OpenStudy (anonymous):

distance = rate times time if the total distance separating them is 50km then our equation will look like: 50 = 53*time + 47*time now all we have to do is solve for time. can you do this?

OpenStudy (anonymous):

How can I solve for t.?

OpenStudy (anonymous):

50=53*t + 47*t we factor out a t on the right side, so 50=t(53+47) does that help?

OpenStudy (anonymous):

i got 100. so what do i do next

OpenStudy (anonymous):

50=t(53+47) 50/(53+47) = t 50/100 = t 5/10 = t 1/2 = t so it is 1/2 and hour till the two drivers are 50km apart